# Proxy Prompt: Endowing SAM & SAM 2 with Auto-Interactive-Prompt for Medical Segmentation

### v1

@@ 4.5 Ablation Studies
Figure 7: The boxes show performance for each $M$ with varying $G$, while in the rightmost plots ($G<M-1$), higher peaks refer to better performance and narrower ranges reflect improved stability. Optimum is achieved when both $M$ and $G$ are maximized.
For each dataset, 16 samples were excluded to ensure a consistent test set across all experiments. To evaluate the impact of training set size $M$, we defined $\mathbf{Train\;set}=\{(\mathbf{I}_{i},\mathbf{M}_{i})\}_{i=1}^{M}$ with $M\in\{2,4,8,16\}$, setting the number of support pairs to $M-1$ during training. For a given $M$, we conducted repeated experiments by randomly selecting multiple independent and diverse $\mathbf{Support\;set}=\{(\mathbf{I}_{i},\mathbf{M}_{i})\}_{i=1}^{G}$ for each inference support size $G\in\{1,2,4,8\}$. Under different values of $M$ and $G$, the averaged dice scores across three datasets and five objects are illuminated in Fig. 7.
The figure shows a positive correlation between segmentation quality and $M$. For each $M$, the dice score monotonically increases with $G$ until $G$ exceeds $M-1$, after which a slight decline may occur. Furthermore, the ridge plot reveals that when $G=1$, the distribution of dice scores is more dispersed due to the highly diverse $\mathbf{Support\;set}$. However, as $G$ increases, the variance in model predictions with various support pairs greatly decreases. For instance, in the top-right plot based on $M=16$, the results of repeated experiments are more concentrated with sharper shape when $G=8$(teal region) compared to $G=1$(charcoal purple region), indicating that the model is less affected by the quality of individual support pairs. In summary, both training dataset size $M$ and inference support size $G$ influence model performance, with $M$ having a notably greater impact. The dice score stabilizes at its highest value when both $M$ and $G$ are maximized, reflecting the optimal performance and stability of our model.
Experiments validating the effectiveness of the two modules are provided in the supplementary materials.
@@ 5 Conclusion
We present the PPG, a pluggable framework designed to endow SAM and SAM 2 with auto-prompting and enhanced their interactive capabilities. Our PPG employs a CSM to extract contextual information from non-target data and a CCM to interpret user intentions, enabling it to generate efficient PP that effectively guides the model output to meet user needs. The PPG also can be implanted into other LVMs, continuously integrating cutting-edge foundation models to meet the evolving demands in widespread medical applications.
Supplementary Material
F Additional Ablation Experiment
We conducted ablation experiments to validate the effectiveness of the Contextual Selective Module (CSM) and Contextual Colorization Module (CCM) modules. To assess the contribution of CCM in interpreting diverse user segmentation intention, we presented results on the FH-PS-AOP (FPA) dataset under two different indications to segment pubic symphysis (PS) and fetal head (FH) . All settings were tested on the same test set, excluding data from 16 subjects for support image-mask pairs. To segment the PS structure in the FPA dataset (FPA-PS), we used the excluded original images and PS masks as inputs to the Proxy Prompt Generator (PPG), replacing the mask with the Fetal Head (FH) mask for FH segmentation (FPA-FH).
Table 2: Ablation experiments on different modules of Proxy Prompt Generator on the FH-PS-AOP dataset, with results reported in terms of Dice scores.
CSM CCM FPA-PS FPA-FH Average
81.5 88.0 84.8
79.4 88.4 83.9
80.3 87.3 83.8
78.3 87.5 82.9
The results in Tab. 2 highlight the contributions of both CSM and CCM to the optimal model performance. When CCM was removed, the segmentation performance for PS structures showed a pronounced decline (2.1% Dice score drop), while FH structures exhibited a slight improvement. This may be attributed to the absence of CCM impairing the PPG’s ability to interpret different user intentions, resulting in markedly divergent segmentation outcomes under varying indications. On the other hand, removing CSM led to an overall degradation in segmentation quality for both FH and PS structures, with the average Dice score decreasing by 1%. When the selective mechanism of CSM was disabled and CCM was simultaneously omitted, model performance fell to the Med-SA baseline. This experiment demonstrates that our model, when enhanced with both modules, achieves optimal and consistent performance across different objects.
G Prompt Strategy
G.1 Settings for Video Dataset.
Considering the high demand for timely operation in real-time image-guided examinations and interventions for physicians, prompt was only required for the first frame of the video data.
1. Box: Minimum bounding rectangle of the target as a bounding box prompt.
2. Point: Center point of the object as a positive point.
3. Mask: Target mask of the patient under examination.
4. Support Image-Mask Pair: Image of the first frame and the corresponding target mask from “non-target” data (i.e., image/video frame of subjects other than the one under examination, such as from retrospective datasets).
G.2 Settings for Image Dataset.
For the models that require prompts in the comparison experiments on image dataset, the following prompt conditions are provided.
1. Point: Since the center point of the disc and vessel mask is not on the target, one random point within the target mask as a positive point prompt.
2. Box: Minimum bounding rectangle of the target as a bounding box prompt.
3. Everything: Automatically segment multiple targets with everything mode. We select the prediction that the highest overlaps with ground truth to calculate the model’s Dice score.
4. Others: The SAMed and AutoSAM models are designed to perform automatic segmentation without manual-given prompt, while we use the support image-mask pair as a prompt.
H Datasets Introduction
Type of image modalities, segmentation objects, and number of samples for the four included datasets are summarized in the table below. The related references are provided in the main text.
Table 3: Dataset summary.
Dataset Modality Segmentation Objects Samples
REFUGE2 Fundus Optic disc and optic cup 2000 images
STARE Fundus Blood vessels in retinal images 20 images
FPA Ultrasound Fetal head and pubic symphysis 4000 images
JNU-IFM Ultrasound Fetal head and pubic symphysis 78 videos

### v2

@@ 4.5 Ablation Studies
We analyze the impact on PP-SAM from multiple aspects, including model architecture, pretrained weights, prompt quality and quantity during inference, and training data size. Experiment settings and details are in the supplementary.
Effectiveness of each model component. The results in Tab. 4 highlight the contributions of both CSM and CCM to the optimal model performance. When CCM was removed, the segmentation performance std increased significantly (2.6% → 3.1%), indicating inconsistencies in multi-object segmentation. Removing CSM led to an overall degradation in segmentation quality, with the average Dice score decreasing by 1.6%. When CSM and CCM were simultaneously omitted, the model achieved the lowest Avg. Dice (84.0%) with the highest std. Dice (3.1%), confirming that CSM and CCM collectively enhance both segmentation accuracy and stability.
Visualization of selective map effectiveness across different support images. Fig. 5 shows that the Selective Map directs the target image to focus on relevant anatomical structures in the support image while suppressing background interference. Across varying support image similarity, from S1 (MSE = 0) to S2 (MSE = 3010.1), it consistently highlights semantically relevant regions, ensuring stable segmentation (Avg. Dice: 95.42%–95.53%).
Effect of encoder architectures in CSM shown in Tab. 11. Vision Transformer achieves better results than ResNet-50 (Avg. Dice: 84.5% vs. 83.9%), which may be attributed to significantly larger number of trainable parameters (89.0M vs. 8.5M). Notably, Vision Mamba attains the highest performance (Dice: 85.9%; IoU: 76.3%) despite having the lowest parameter count (7.9M).
Encoder in CSM Turnable Param (M) Avg. Dice $\uparrow$ (%) Avg. IoU $\uparrow$ (%)
Vision Mamba 7.9 85.9 76.3
Vision Transformer 89.0 84.5 74.0
ResNet-50 8.5 83.9 72.6
Table 5: Ablation on different encoder architectures in CSM.
Ablation on pretrained foundation models. Tab. 6 shows that our Proxy Prompt consistently enhances SAM-based models, including MedSAM, which was pretrained on extensive medical data. Surprisingly, MedSAM with our prompt underperforms compared to SAM with the same prompt, likely due to its strong reliance on Box prompts during pretraining, limiting the adaptability to our prompt. Deep analysis is in the supplementary.
Method Pretrained Weight Prompt Method Avg. Dice $\uparrow$ (%)
MedSAM MedSAM-ViT-B Box 74.5
Proxy Prompt 81.3 (+6.8)
SAM SAM-ViT-B Box 56.1
Proxy Prompt 85.1 (+29.0)
SAM-ViT-H Box 60.3
Proxy Prompt 85.9 (+25.6)
Table 6: Performance with different pretrained foundation models.
Influence of prompt quality via support mask variations. We evaluate six levels of support mask quality, recording each mask’s Dice score as Prompt Dice in Fig. 9. The results shows segmentation improves (visualized by intensifying red shades) with prompt quality, dropping sharply only in $Empty$ level. Our model remains robust regardless of whether masks come from normal users or senior doctors. Line chart is in the supplementary.
Empty Low Medium High SD NU
Prompt Dice (%) 00.0 66.6 81.6 100 88.3 74.8
Output Dice (%) 47.9 77.8 86.5 86.8 86.6 86.5
Table 7: Model performance across prompt qualities, including senior doctor (SD) and normal user (NU). Prompt Dice reflects support mask quality, while Output Dice measures performance.
@@
Figure 6: The boxes show performance for each $M$ with varying $K$, while in the rightmost plots ($G<M-1$), higher peaks refer to better performance and narrower ranges reflect improved stability. Optimum is achieved when both $M$ and $K$ are maximized.
Impact of training data size $M$ and support pair quantity $K$ in inference. Fig. 11 shows that performance improves with $M$. For each $M$, Dice increases with $K$ until $K$ exceeds $M-1$, after which a slight decline may occur. The $M=16$ ridge plot indicates that as $K$ increases, prediction variance decreases significantly. Overall, both training dataset size $M$ and support size $K$ influence model performance, which stabilizes at its peak when both are maximized. Experiment details are in the supplementary.
@@ 5 Conclusion
We present the PPG, a pluggable framework designed to endow SAM and SAM 2 with auto-prompting and enhanced their interactive capabilities. Our PPG employs a CSM to extract contextual information from non-target data and a CCM to interpret user intentions, enabling it to generate efficient PP that effectively guides the model output to meet user needs.